12:42 UT We had more problems so far this month involving mainly Ant 14, perhaps exacerbated by more snow. The X polarization on the high-frequency receiver showed a 15 dB loss of RF power, so Mark replaced the cable on Dec. 2. However, he accidentally left a glove at the focus, which fell down into the RX Select gear mechanism, so it was stuck at some intermediate position. The low-frequency receiver was still getting some signal, but no calibration was possible for the high-frequency receiver. Yesterday Mark realized the problem and managed to get the receiver unstuck and working again. Last night we finally got some reasonable calibration, which showed that a 5 ns delay difference was introduced. I updated the delay so hopefully starting with today's data the observations will again be nominal. Despite these problems, and the lack of a good calibration for about a week, our daily pipeline images still look pretty good!
14:25 UT Just a note to say today's pipeline output, the full disk images from Dec. 9, look very good--some of the best I have seen. These are possibly due to our finally getting a good calibration for that date. The main remaining problem is Ant 13, whose calibration data are of very low quality except at the lowest frequencies. I have not had time to figure out the cause.